OpenStudy (moonlitfate):

Find the derivative of the function: y = arctan(2x^2-3)

OpenStudy (moonlitfate):

I have most of the problem done already, it's just one part that's messing me up. The algebra ironically, lol.

OpenStudy (amistre64):

solve by first principles, or by takinga tangent of both sides?

OpenStudy (moonlitfate):

I took the tangent of both sides.

OpenStudy (math&ing001):

\[(\arctan(u))'=\frac{ u' }{ 1+u ^{2} }\]

OpenStudy (math&ing001):

Take u=2x²-3

OpenStudy (amistre64):

tan y =f(x) sec^2(y) y' = f' y' = f'/(tan^2(y) + 1)

OpenStudy (amistre64):

but tan(y) = f(x) sooo y' = f(x)/([f(x)]^2+1)

OpenStudy (amistre64):

lost a ' on top lol

OpenStudy (moonlitfate):

\[u = 2x-3\]\[u' = 2\]

OpenStudy (amistre64):

so that hard part you say is: 4x ------------- (2x^2-3)^2 + 1

OpenStudy (math&ing001):

Yeah, and replace on the formula.

OpenStudy (math&ing001):

It's 2x^2-3 btw.

OpenStudy (moonlitfate):

Yeahm like I went and plug that in and got: \[y' = \frac{ 2 }{ 1+(2x^2-3)^2 }\]

OpenStudy (moonlitfate):

Unless I did something wrong there.

OpenStudy (math&ing001):

Should be 4x on top.

OpenStudy (moonlitfate):

... I left off the the exponent of u... :/

OpenStudy (moonlitfate):

Let me to re-do that, actually... one sec. And it'll help mwe with learning this.

OpenStudy (math&ing001):

Ok take your time.

OpenStudy (moonlitfate):

@math&ing001 -- But, yes, that is where I'm stuck. \[y' =\frac{ 4x }{ 1+(2x^2-3)^2 }\]

OpenStudy (math&ing001):

You need to simplify it ?

OpenStudy (moonlitfate):

Yeah.

OpenStudy (math&ing001):

I'd start with the term (2x^2-3)^2

OpenStudy (moonlitfate):

That would just be using the FOIL method? My Algebra skills are just rusty, not to mention Calculus. But, I'm willing to work through this, haha

OpenStudy (math&ing001):

You could use FOIL or the formula for binomials : (a+b)^2 = a^2 + 2ab + b^2

OpenStudy (moonlitfate):

Ahhh, okay. I had used FOIL and got: \[(2x^2-3)^2 = 4x^4-12x^2+9\]

OpenStudy (math&ing001):

Yep that's correct !

OpenStudy (math&ing001):

@MoonlitFate

OpenStudy (math&ing001):

So that would give us: \[y'=\frac{ 4x }{ 4x ^{4}-12x ^{2}+10 }\] I don't think it can be further simplified.

OpenStudy (moonlitfate):

Oh, really? Maybe that's why... I'm trying to figure out an answer out, and keep coming up with nothing.